OpenStudy (anonymous):

Help. I need somebody. Help not just anybody. Help. g+4-3g=1+g

OpenStudy (anonymous):

\[g+4-3g=1+g\] combine like terms first on the left, get \[-2g+4=1+g\]

OpenStudy (anonymous):

then add \(2g\) to both sides get \[4=1+3g\]

OpenStudy (anonymous):

subtract \(1\) gives you \[3=3g\] making \(g=1\)

OpenStudy (wolf1728):

Amuru - seems you are a Beatles fan?

OpenStudy (anonymous):

Wh-? How do I seem like a Beatles fan? I'm chill with them but not a huge fan. And thanks satellite73

OpenStudy (anonymous):

yw it was the Help reference